OpenStudy (anonymous):

describe how to transform (5 sqrt x^7)^3 into an expression with a rational exponent .

OpenStudy (anonymous):

@DeadShot @ash2326 @mathstudent55

OpenStudy (anonymous):

@Hero Hey I wa wondering if you could help a lil bit

hero (hero):

Hint: \(\sqrt{x} = x^{1/2}\)

OpenStudy (mathstudent55):

In general, \( \Large x ^{\frac{m}{n}} = \sqrt[n]{x^m} = (\sqrt[n]{x})^m \)

hero (hero):

Just to verify, is it\(\left(5 \sqrt{x^7}\right)^3\) or \(\left(\sqrt[5]{x^7}\right)^3\)

hero (hero):

@crystelle

OpenStudy (anonymous):

Oh wow. I exit it and now they changed the questions. Same form but its now (\[(^3 \sqrt{x}^4)^5 \]

OpenStudy (anonymous):

@hero @mathstudent55